OpenStudy (anonymous):

help mee please. If x-y=20, and x^2-y^2=50, what is the value of X

OpenStudy (mertsj):

Replace y in the second equation with x-20 from the first equation.

OpenStudy (loser66):

x-y not x-20 Mertsj

OpenStudy (mertsj):

How so?

OpenStudy (anonymous):

@mertsj, are you using the elimination by substitution method

OpenStudy (mertsj):

When you replace a variable, that means substitute something else for it.

OpenStudy (anonymous):

ok. so now its x^2 - (x-20)^2=50

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so after you distribute the negative sign it becomes x^2(-x+20)^2=50

OpenStudy (mertsj):

\[x^2-(x^2-40x+400)=50\]

OpenStudy (mertsj):

\[x^2-x^2+40x-400=50\]

OpenStudy (anonymous):

what did you just do.

OpenStudy (mertsj):

\[40x=450\]

OpenStudy (mertsj):

Expand (x-20)^2 Remove the parentheses by changing everything inside to its opposite.

OpenStudy (mertsj):